what is 5(d+4) = 7(d-2)
Answer by tommyt3rd(5050)

I assume you mean to solve for d...

5(d+4)=7(d-2)

5d+20=7d-14


5d-5d+20+14=7d-5d-14+14


34=2d

d=17


:)